1. A pill weighing 325mg contains Captopril, a medication for high blood pressure. What is the concentration of Captopril in ppm?
  2. A patient is given 1000mL of 5 percent glucose C6H12O6 solution intravenously in a 12 hour period. How many grams of glucose did the patient receive per hour?
1
Expert's answer
2022-02-15T14:39:04-0500
  1. The concentration of Captopril in ppm depends on its total content in the pill.
  2. % = (100 x msolute) / (msolvent + msolute)

5% = (100 x msolute) / 1000

1000 ml = 1000 g

msolute (C6H12O6) = (5 x 1000) / 100 = 50 g

msolute (C6H12O6) per 1 hour = 50 / 12 = 4.2 g